it is necessary to have a 70% antifreeze solution in a radiator of a certain car. the radiator now has 20 ml of 60% solution. how much of this solution should be drained and replaced with 100% antifreeze to get the desired strength
Answer by lwsshak3(11628)

it is necessary to have a 70% antifreeze solution in a radiator of a certain car. the radiator now has 20 ml of 60% solution. how much of this solution should be drained and replaced with 100% antifreeze to get the desired strength.
**
x=ml of 60% solution to be drained and replaced with 100% antifreeze
20-x=ml of 60% solution remaining
..
60%(20-x)+100%x=70%*20
12-0.6x+x=14
0.4x=2
x=5 ml
ans:
ml of 60% solution to be drained and replaced with 100% antifreeze=5